# Theoretical Studies of Sub-THz Active Split-Ring Resonators for Near-Field Imaging

### v1

@@ VI Conclusion
In this paper, the analysis and design of ASRRs were presented. Two key features, namely the tunable boosted quality factor and the switchability, enable the dense integration of ASRRs in scalable 2D arrays. The near fields generated by these arrays are employed for high-resolution imaging of biological tissues, which find applications in disease diagnostics. The resonators are highly sensitive to both the real and imaginary parts of the sample permittivity and also exhibit an extra gain equal to the square of the $Q$-boosting ratio to the changes in dielectric losses. Equivalent LC models for the ASRRs were developed and shown to closely match the behavior of the HFSS EM models. The optimum coupling conditions and pixel saturation were studied. The nonlinear effects of the $g_{m}$ circuits inside the resonator and their impact on de-Qing the resonator, when driven by excessive power, were investigated, leading to an optimum input power. The effects of various noise sources, including ASRR device white and flicker noise, the supply noise, and the input phase noise, on the output phase noise were analyzed. Ultimately, these studies culminated in providing design guidelines for maximizing the detection SNR while minimizing the power consumption.
@@ Appendix B Nonlinear Transconductance Calculations
Since a sine wave spends most of its time at its peaks, a sharp drop in $g_{m,avg}$ is expected when the voltage swing enters the triode region. Carrying out the integrations in (63), $g_{m,avg}$ can be derived from the following expression

### v2

@@ VI Conclusion
In this paper, the analysis and design of ASRRs were presented. Two key features, namely the tunable boosted quality factor and the switchability, enable the dense integration of ASRRs in scalable 2D arrays. The near fields generated by these arrays are employed for high-resolution imaging of biological tissues, which find applications in disease diagnostics. The resonators are highly sensitive to both the real and imaginary parts of the sample permittivity and also exhibit an extra gain equal to the square of the $Q$-boosting ratio to the changes in dielectric losses. Equivalent LC models for the ASRRs were developed and shown to closely match the behavior of the HFSS EM models. The optimum coupling conditions and pixel saturation were studied. The nonlinear effects of the $g_{m}$ circuits inside the resonator and their impact on de-Qing the resonator, when driven by excessive power, were investigated, leading to an optimum input power. The effects of various noise sources, including ASRR device white and flicker noise, the supply noise, and the input phase noise, on the output phase noise were analyzed. Ultimately, these studies culminated in providing design guidelines for maximizing the SNR while minimizing the power consumption.
@@ Appendix B Nonlinear Transconductance Calculations
A sine wave spends most of its time at its peaks, resulting in a sharp drop in $g_{m,avg}$ when the voltage swing enters the triode region. Carrying out the integrations in (63), $g_{m,avg}$ can be derived from the following expression